If 180.3g of aluminum reacts with 103.4g of sulfuric acid then 203.7 grams of Aluminum Sulfate and _______g of Hydrogen gas would be produced.
How many grams of Hydrogen gas is produced in this instance?
1
Expert's answer
2019-09-26T05:54:43-0400

First, the limiting reactant of the reaction must be determined. For the purpose, number of moles (n) of reactants must be calculated.

According to the reaction:

2Al + 3 H2SO4 --> Al2(SO4)3 + 3H2

n (Al) = 180.3 g / (2 × 27 g/mol) = 3.33 mol

n (H2SO4) = 103.4 g / (3 × 98 g/mol) = 0.35 mol

As  sulfuric acid is a limiting reactant, the mass of hydrogen gas is:

m (H2) = 103.4 g × 6 g/mol / 294 g/mol = 2.11 g


Answer: 2.11 g of hydrogen gas.